# Quantum cohomology of projective bundles

### v3

@@ 5.4. Fourier transform of the $J$-function
Remark 5.7.
The leading term $-c_{1}(V)\log\lambda_{j}$ of $\sigma_{j}$ accounts for difference of the degree of $Q^{d}$ between $B$ and $V$.
@@
Computing the formal stationary phase asymptotics of (5.4) at $Q=0$ over the Artinian local ring $H^{*}(B)$, we see that $r\lambda_{j}+\sigma_{j}|_{Q=0}$ is given by the critical value of (cf. (3.7))
@@
at a unique $H^{*}(B)$-valued critical point $\hat{\lambda}_{j}\in\lambda_{j}(1+\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1},\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!])$, which is a solution $\lambda$ to the equation $q=e_{\lambda}(V)e^{-\tau^{\prime}(\lambda)}$. The critical point $\hat{\lambda}_{j}$ and the critical value $f(\hat{\lambda}_{j})$ are the eigenvalues of the quantum products $(p\star)|_{Q=0}$ and $(E_{\mathbb{P}(V)}\star)|_{Q=0}$ in the $H^{*}(B)$-algebra $QH^{*}(\mathbb{P}(V))|_{Q=0}$ respectively. (It is easy to see that $H^{*}(B)$ acts by cup product on $QH^{*}(\mathbb{P}(V))|_{Q=0}$.)
@@ 5.5. Projections induced by the Fourier transformations
Let $\sigma_{j}$ be as in Proposition 5.6. We regard it as a multi-valued map from $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ to $H^{*}(B)$, where $r^{\prime}$ is as in (5.1), and consider the pull-back $\sigma_{j}^{*}\operatorname{QDM}(B)$ of the quantum $D$-module of $B$ by $\sigma_{j}$. We define $\sigma_{j}^{*}\operatorname{QDM}(B)$ to be the module
$\sigma_{j}^{*}\operatorname{QDM}(B)=H^{*}(B)[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
equipped with the pull-back $\sigma_{j}^{*}\nabla$ of the quantum connection of $B$:
@@
$(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(q\partial_{q}\sigma_{j})\star_{\sigma_{j}}$
@@
$(\sigma_{j}^{*}\nabla)_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}E_{B}\star_{\sigma_{j}}+\mu_{B}$
and the pairing $P_{B}$ as in (2.8). The Divisor Equation show that these pull-backs are well-defined and define maps
$q^{-k/r}(\sigma_{j}^{*}\nabla)_{\tau^{i,k}},\ (\sigma_{j}^{*}\nabla)_{q\partial_{q}},\ (\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}},\ (\sigma_{j}^{*}\nabla)_{z\partial_{z}}\colon\sigma_{j}^{*}\operatorname{QDM}(B)\to z^{-1}\sigma_{j}^{*}\operatorname{QDM}(B).$
The operator $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ has poles along the infinity disivor $(q^{-1/r^{\prime}}=0)$, but this is an apparent singularity: $\partial_{\tau^{i,k}}$ viewed as a vector field on $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ has poles of the same order along $(q^{-1/r^{\prime}}=0)$.
We introduce a localized version along the infinity disivor $(q^{-1/r^{\prime}}=0)$ as follows:
$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ $:=\sigma_{j}^{*}\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$
@@
The Fourier transformation $\mathscr{F}_{j}$ induces a map $\Pi_{j}\colon\operatorname{QDM}_{S^{1}}(V)\to\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules such that
(5.5) $\mathscr{F}_{j}(M_{V}(\tau)s)=M_{B}(\sigma_{j})\Pi_{j}(s)\qquad\text{for all $s\in\operatorname{QDM}_{S^{1}}(V)$.}$
The map $\Pi_{j}$ satisfies the following:
@@
    $\Pi_{j}$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(B)$ and the degree of the variables as in Proposition 5.6;
@@
By differentiating the identity $\mathscr{F}_{j}(M_{V}(\tau)1)=M_{B}(\sigma_{j})\upsilon_{j}$ from Proposition 5.6 by $\tau^{i,k}$ and using the differential equations (2.7) for $M_{V}$ and $M_{B}$, we obtain
$\mathscr{F}_{j}(M_{V}(\tau)\phi_{i}\lambda^{k})=M_{B}(\sigma_{j})z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}.$
Therefore the $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-module map $\Pi_{j}$ sending $\phi_{i}\lambda^{k}$ to $z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}$ satisfies (5.5). Parts (1)-(5) follow immediately from (5.5), Proposition 5.5 and the intertwining properties (2.7), (2.21) of $M_{V},M_{B}$. Part (6) follows from the homogeneity of $\sigma_{j}+c_{1}(V)\log\lambda_{j}$, $\upsilon_{j}$ in Proposition 5.6. Part (7) follows by a direct computation of
@@ 5.6. End of the proof
$\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}.$
By Theorem 4.2 and Proposition 5.9, this is a homomorphism of $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$-modules homogeneous of degree $-(r-1)$, and intertwines the connection $\hat{\tau}^{*}\nabla$ with
@@
We note that the pull-back $(\sigma_{j}+r\lambda_{j})^{*}\nabla$ of the quantum connection of $B$ by $\sigma_{j}+r\lambda_{j}$ is well-defined [^8] Although $M_{B}(\sigma_{j}+r\lambda_{j})$ is ill-defined in the $q^{-1/r}$-adic topology. and that
@@
(5.6) $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm loc}$
that is fully compatible with the connection. We now restrict the parameter $\tau$ to lie in $\mathsf{H}:=\bigoplus_{k=0}^{r-1}H^{*}(B)\lambda^{k}\subset H^{*}_{S^{1}}(V)$: the mirror map $\tau\mapsto\hat{\tau}(\tau)$ in Theorem 4.2 restricted to $\mathsf{H}$ gives an isomorphism between $\mathsf{H}$ and $H^{*}(\mathbb{P}(V))$ in the formal neighbourhood of $q=Q=0$ since $\hat{\tau}(\tau)|_{q=Q=0}=\kappa(\tau)$. We compose the map $\sigma_{j}$ with the inverse $\hat{\tau}\mapsto\tau=\tau(\hat{\tau})$ of this isomorphism and define
@@
Because of the homogeneity of $\hat{\tau}(\tau)$ from Theorem 4.2(6), the inverse mirror map $\tau(\hat{\tau})$ belongs to $H^{*}(B)[q][\![Q,\hat{\tau}]\!]$. Therefore the composition is well-defined; $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ belongs to $H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$ and is homogeneous of degree $2$. Moreover we have
(5.7) $\varsigma_{j}(\hat{\tau})\Bigr{|}_{Q=0}\in r\lambda_{j}-c_{1}(V)\log\lambda_{j}+\alpha+\sum_{k=0}^{r-1}\sum_{i=0}^{s}\hat{\tau}^{i,k}\phi_{i}\lambda_{j}^{k}(1+\beta_{k})+(\hat{\tau})^{2}H^{*}(B)$
for some $\alpha,\beta_{k}\in\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1}]\!]$, where $(\hat{\tau})\subset\mathbb{C}(\!(q^{-1/r})\!)[\![\hat{\tau}]\!]$ denote the ideal generated by $\hat{\tau}^{i,k}$ with $0\leq i\leq s$, $10\leq k\leq r-$ 1. This follows by Theorem 4.2(7) and Proposition 5.6. The above map induces a $\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$-module map
$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
compatible with the connection and homogeneous of degree $-(r-1)$. Combining Theorem 4.2(8) and Proposition 5.9(7), we have
@@
for $10\leq k\leq r-$ 1 and therefore $\Phi$ induces an isomorphism over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$. Inverting $\Phi|_{Q=\hat{\tau}=0}$, we obtain
@@
$(\varsigma_{j}^{*}\nabla)_{q\partial_{q}}\Bigr{|}_{Q=\hat{\tau}=0}=q\partial_{q}+\frac{1}{z}\left(\lambda_{j}+\frac{c_{1}(V)}{r}+\alpha^{\prime}\right)$
where $c_{1}(V)$ and $\alpha^{\prime}=q\partial_{q}\alpha\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$ acts by cup product. Writing $x=q^{-1/r}$, we find that $v(x):=P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=v_{0}+v_{1}x+v_{2}x^{2}+\cdots$ satisfies a differential equation of the form
@@
where $S,N_{i}$ are mutually commuting endomorphisms on $\operatorname{End}(\bigoplus_{j=0}^{r-1}H^{*}(B)))$ such that $S$ is semisimple and $N_{0}$ is nilpotent. We obtain a sequence of equations
@@
From $Sv_{0}=N_{i}v_{0}=0$ and (*$*_{0}$) we have $Sv_{1}=0$. Then (*$*_{1}$) implies $S^{2}v_{2}=0$. Since $S$ is semisimple, $Sv_{2}=0$ and (*$*_{1}$) implies $(1+N_{0})v_{1}=0$; the nilpotence of $N_{0}$ implies $v_{1}=0$. In this way we can show inductively that $v_{1}=v_{2}=\cdots=0$. Therefore $P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=I$; in other words,
@@
We now use the fact that $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ is flat with respect to the connection on the endomorphism bundle of $\operatorname{QDM}(\mathbb{P}(V))$. It satisfies a differential equation in $(Q,\hat{\tau})$ with at most logarithmic singularities at $Q=0$ and nilpotent residues. Expanding $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ in power series in $Q$ and $\hat{\tau}$ and using a similar argument as before, we can easily show that $P_{1}^{-1}\Phi^{T}P_{2}\Phi=I$ identically. Thus $\Phi$ preserves the pairing.
@@
The map (5.6) glues the connections $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$; $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ is defined around $q=Q=\boldsymbol{\tau}=0$ and $(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$ is defined around $q^{-1/r^{\prime}}=q^{-c_{1}(V)/r}Q=q^{\bullet/r}\boldsymbol{\tau}=0$. They glue to a connection over a global “Kähler moduli space” obtained by adding the infinity divisor $(q=\infty)$ to $\operatorname{Spf}(\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!])$. See Figure 1 for the global Kähler moduli space.

### v4

@@ 5.4. Fourier transform of the $J$-function
Computing the formal stationary phase asymptotics of (5.6) at $Q=0$ over the Artinian local ring $H^{*}(B)$, we see that $r\lambda_{j}+\sigma_{j}|_{Q=0}-\frac{1}{r}c_{1}(V)\log q$ is given by the critical value of (cf. (3.4))
@@
at a unique $H^{*}(B)$-valued critical point $\hat{\lambda}_{j}\in\lambda_{j}(1+\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1},\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!])$, which is a solution $\lambda$ to the equation $q=e_{\lambda}(V)e^{-\tau^{\prime}(\lambda)}$. Here $\hat{\lambda}_{j}$ is an eigenvalue of the quantum product $(p\star)|_{Q=0}$ in the $H^{*}(B)$-algebra $QH^{*}(\mathbb{P}(V))|_{Q=0}$. (It is easy to see that $H^{*}(B)$ acts by cup product on $QH^{*}(\mathbb{P}(V))|_{Q=0}$.)
@@ 5.5. Projections induced by the Fourier transformations
Let $\sigma_{j}$ be as in Proposition 5.6. We regard it as a map from $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ to $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,\sigma]\!]$, where $r^{\prime}$ is as in (5.1), and consider the pull-back of the extended quantum $D$-module $\operatorname{QDM}(B)_{\rm ext}$ (5.3) of $B$ by $\sigma_{j}$. The pull-back is defined to be the module
$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}:=H^{*}(B)[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
equipped with the pull-back $\sigma_{j}^{*}\nabla$ of the connection for $\operatorname{QDM}(B)_{\rm ext}$:
@@
$(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(-(c_{1}(V)/r)\star_{\sigma_{j}}+(q\partial_{q}\sigma_{j})\star_{\sigma_{j}})$
@@
$(\sigma_{j}^{*}\nabla)_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\sigma_{j}})+\mu_{B}$
and the pairing $P_{B}$ as in (2.8). The Divisor Equation shows that these pull-backs are well-defined and define maps
$q^{-k/r}(\sigma_{j}^{*}\nabla)_{\tau^{i,k}},\ (\sigma_{j}^{*}\nabla)_{q\partial_{q}},\ (\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}},\ (\sigma_{j}^{*}\nabla)_{z\partial_{z}}\colon\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}\to z^{-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}.$
The operator $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ has poles along the infinity divisor $(q^{-1/r^{\prime}}=0)$, but this is an apparent singularity: $\partial_{\tau^{i,k}}$ viewed as a vector field on $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ has poles of the same order along $(q^{-1/r^{\prime}}=0)$.
Similarly to (5.3), we denote by $\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ the restriction of $\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}$ to the localized base $\operatorname{Spf}\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$.
$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ $:=\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}\otimes_{\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$
@@
The Fourier transformation $\mathscr{F}_{j}$ induces a map $\Pi_{j}\colon\operatorname{QDM}_{S^{1}}(V)\to\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules such that
(5.7) $\mathscr{F}_{j}(M_{V}(\tau)s)=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\Pi_{j}(s)$
for all $s\in\operatorname{QDM}_{S^{1}}(V)$. The map $\Pi_{j}$ satisfies the following:
@@
    $\Pi_{j}$ is homogeneous of degree (r1)-(r-1) with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(B)$ and the degree of the variables as in Proposition 5.6;
@@
By differentiating the identity $\mathscr{F}_{j}(M_{V}(\tau)1)=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$ from Proposition 5.6 by $\tau^{i,k}$ and using the differential equations (2.7) for $M_{V}$ and $M_{B}$, we obtain
$\mathscr{F}_{j}(M_{V}(\tau)\phi_{i}\lambda^{k})=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}.$
Therefore the $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-module map $\Pi_{j}$ sending $\phi_{i}\lambda^{k}$ to $z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}$ satisfies (5.7). Parts (1)-(5) follow immediately from (5.7), Proposition 5.5 and the intertwining properties (2.7), (2.21) of $M_{V},M_{B}$. Part (6) follows from the homogeneity of $\sigma_{j}$, $\upsilon_{j}$ in Proposition 5.6. Part (7) follows by a direct computation of
@@ 5.6. End of the proof
$\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}.$
By Theorem 4.2 and Proposition 5.8, this is a homomorphism of $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$-modules homogeneous of degree (r1)-(r-1), and intertwines the connection $\hat{\tau}^{*}\nabla$ with
@@
We note that the pull-back $(\sigma_{j}+r\lambda_{j})^{*}\nabla$ of the quantum connection of $B$ by $\sigma_{j}+r\lambda_{j}$ is well-defined [^10] Despite the fact that the pull-back $M_{B}(\sigma_{j}+r\lambda_{j})$ of the fundamental solution is ill-defined in the $q^{-1/r^{\prime}}$-adic topology. and that
@@
(5.8) $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
[…]
@@
[…]
(5.9) $\varsigma_{j}(\hat{\tau})\Bigr|_{Q=0}\in r\lambda_{j}-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\alpha_{j}+\sum_{k=0}^{r-1}\sum_{i=0}^{s}\hat{\tau}^{i,k}(\phi_{i}+\beta_{j,i,k})\lambda_{j}^{k}+(\hat{\tau})^{2}H^{*}(B)$
for some $\alpha_{j},\beta_{j,i,k}\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$, where $(\hat{\tau})\subset\mathbb{C}(\!(q^{-1/r})\!)[\![\hat{\tau}$]\!] denote the ideal generated by $\hat{\tau}^{i,k}$ with $0\leq i\leq s$, $0\leq k\leq r-1$. Parts (4) and (5) of Theorem 5.1 follow from (5.9). The above map induces a $\mathbb{C}[z,q][\![Q,\hat{\tau}$]\!]-module map
$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
compatible with the connection and homogeneous of degree (r1)-(r-1). Combining Theorem 4.2(8) and Proposition 5.8(7), we have
@@
for $0\leq k\leq r-1$ and therefore $\Phi$ induces an isomorphism over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!]. This is part (6) of Theorem 5.1. Inverting $\Phi|_{Q=\hat{\tau}=0}$, we obtain
@@
$(\varsigma_{j}^{*}\nabla)_{q\partial_{q}}\Bigr|_{Q=\hat{\tau}=0}=q\partial_{q}+\frac{1}{z}\left(\lambda_{j}-\frac{c_{1}(V)}{r}+q\partial_{q}\alpha_{j}\right)$
where $c_{1}(V)$ and $q\partial_{q}\alpha_{j}\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$ acts by cup product. Writing $x=q^{-1/r}$, we find that $v(x):=P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=v_{0}+v_{1}x+v_{2}x^{2}+\cdots$ satisfies a differential equation of the form
@@
where $S,N_{i}$ are mutually commuting endomorphisms on $\operatorname{End}(\bigoplus_{j=0}^{r-1}H^{*}(B))$ such that $S$ is semisimple and $N_{0}$ is nilpotent. We obtain a sequence of equations
@@
From $Sv_{0}=N_{i}v_{0}=0$ and ($_{0}$) we have $Sv_{1}=0$. Then ($_{1}$) implies $S^{2}v_{2}=0$. Since $S$ is semisimple, $Sv_{2}=0$ and ($_{1}$) implies $(1+N_{0})v_{1}=0$; the nilpotence of $N_{0}$ implies $v_{1}=0$. In this way we can show inductively that $v_{1}=v_{2}=\cdots=0$. Therefore $P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=I$; in other words,
@@
[…]
@@
The map (5.8) glues the connections $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext}$; $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ is defined around $q=Q=\boldsymbol{\tau}=0$ and $(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext}$ is defined around $q^{-1/r^{\prime}}=q^{-c_{1}(V)/r}Q=q^{\bullet/r}\boldsymbol{\tau}=0$. They glue to a connection over a global “Kähler moduli space” obtained by adding the infinity divisor $(q=\infty)$ to $\operatorname{Spf}(\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!])$. See Figure 1.